OpenStudy (anonymous):

Find the vertex, value of p, axis of symmetry, focus, and directrix of the parabola. x=1/24(y−4)^2

OpenStudy (anonymous):

x=1/24(y-4)^2

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

\[4p(x-h)=(y-k)^2 \text{ is the form of the parabola we want to compare to } \\ x=\frac{1}{24}(y-4)^2\] (h,k) is vertex focus is (h+p,k)

myininaya (myininaya):

 parabola you have there looks something like that